OpenStudy (anonymous):

5over x =10 over 3x + 4

OpenStudy (anonymous):

what do you mean by "over"?

OpenStudy (anonymous):

Like divided by?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I dont know how to type in a fraction

OpenStudy (anonymous):

so could u draw what it looks like?

OpenStudy (anonymous):

(5/x)=(10/3x)+4 (multiply both sides by 3x) 15=10+4(3x) 15-10=12x 5=12x x=(5/12)

OpenStudy (anonymous):

5/x =10/3x + 4

OpenStudy (anonymous):

why did you place the 10 with the 4?

OpenStudy (anonymous):

I did not ...

OpenStudy (anonymous):

If you multiply both sides by 3x (to get rid of the fraction on the right hand side) you get 3x(5/x)= 3x(10/3x) + 4(3x)

OpenStudy (anonymous):

3*5=10 + 12x 15=10+12x

OpenStudy (anonymous):

take the 10 to the other side 15-10=12x 5=12x (divide by 12 on both sides) (5/12)=x

OpenStudy (anonymous):

Is it clear?